OpenStudy (anonymous):

Show all work to simplify (2/x)+(2/x+1) - (2/x+2) Use complete sentences to explain how to simplify this

OpenStudy (anonymous):

Remember to list all restrictions.

jhonyy9 (jhonyy9):

x not can being 0,-1,-2 2/x + (2/x+1) -(2/x+2)=(2(x+1)(x+2))+(2(x(x+2))-(2(x(x+1))/(x(x+1)(x+2))= =((2x+2)(x+2)+(2x2+4x)-2x2-2x)/(x(x+1)(x+2))= =(2x2 +6x +4+2x2 +4x -2x2 -2x)/(x(x+1)(x+2)= =(2x2 +8x +4)/(x(x+1)(x+2))